# Preserving Angles Improves Feature Distillation of Foundation Models

### v2

@@ Abstract
Knowledge distillation approaches compress models by training a student network using the classification outputs of a high quality teacher model, but can fail to effectively transfer the properties of computer vision foundation models from the teacher to the student. While it has been recently shown that feature distillation—where a teacher model’s output features are replicated instead—can reproduce performance for foundation models across numerous downstream tasks, they fall short in matching critical properties such as robustness and out-of-distribution (OOD) detection performance. This paper overcomes this shortcoming by introducing Cosine-similarity Preserving Compression (CosPress), a feature distillation technique that learns a mapping to compress the latent space of the teacher model into the smaller latent space of the student, by preserving the cosine similarities between image embeddings. This enables direct optimisation of the student network and produces a more faithful reproduction of the teacher’s properties. It is shown that distillation with CosPress on a variety of datasets, including ImageNet, produces more accurate models with greater performance on generalisability, robustness and OOD detection benchmarks, and that this technique provides a competitive pathway for training highly performant lightweight models on small datasets. Code is available at https://github.com/emannix/cospress.
Figure 1: Effectiveness of feature distillation. Comparison of Proteus and CosPress on a range of benchmarks for distilling features from a DINOv2 ViT-B/14 model into a smaller ViT-S/14 network on ImageNet-1K.
@@
1 Introduction
Deep learning computer vision approaches have become the standard approach for automating vision problems across a range of fields, from medical imaging [55] and analysis of satellite imagery [4] to detecting weapons in luggage [1]. However, models trained with commonly used supervised learning approaches can have poor robustness [3, 16] and struggle to detect out-of-distribution (OOD) data [48, 30].
By leveraging large Vision Transformer (ViT) architectures and pretraining on large and diverse datasets, foundation models in computer vision comprise a significant step forward toward adressing to these challenges, providing significantly improved generalisation ability and robustness in comparison to purely supervised approaches [32, 34]. Large ViT models enjoy superior performance after pre-training [53], and can be distilled to produce smaller models that are more practical for deployment. For example, smaller DINOv2 foundation models were distilled from their largest variant with 1.1 billion parameters by optimizing the self-supervised training objective with a frozen teacher [32]. This approach is not replicable, as it was conducted on the proprietary LVD-142M dataset and the teacher head weights were never publicly released.
Nevertheless, knowledge distillation approaches [17] that leverage the classification outputs of the DINOv2 models trained on a particular datasets can be used to train performant. However, it has been shown that fundamental properties of the foundation model need not transfer to the student models, impacting generalisation performance on downstream tasks [56]. Feature distillation approaches that use a Mean Squared Error (MSE) or $L_{2}$ loss and a student head to map the activations in the latent space of the student to the teacher are more effective in this respect. The recent Proteus [56] approach has shown that it is possible to distill the DINOv2 models on ImageNet-1K and obtain comparable performance on downstream classification and segmentation tasks. However, sub-optimal results are still obtained for key robustness metrics, as well as generalisability and performance in dense tasks such as image segmentation. Most concerning, however, is that models distilled using Proteus do not produce faithful reproductions of the latent space of the teacher model, as shown by their severely reduced performance on OOD detection tasks (Tab. 1).
@@ 1 Introduction
Proteus ViT-Ti/14 ViT-S/14 64.17 85.73 74.22 67.97
CosPress ViT-Ti/14 ViT-S/14 70.49 77.29 91.03 37.21
ViT-S/14 72.58 74.12 92.67 29.55
Proteus ViT-S/14 ViT-B/14 61.19 94.56 61.92 86.78
CosPress ViT-S/14 ViT-B/14 73.5 73.84 92.93 28.98
This paper presents Cosine-similarity Preserving Compression (CosPress), a feature distillation approach that addresses such shortcomings by learning a teacher head mapping that compresses the latent space of the teacher model into the latent space of the student. CosPress achieves this by preserving cosine similarities between points in the teacher’s latent space and allows the student to be directly optimised. This significantly improves the faithfulness of the learned student on OOD detection (Tab. 1) and robustness benchmarks in comparison to Proteus, while achieving competitive performance across all of the considered challenges—including classification accuracy, generalisation and semantic segmentation (Fig. 1). CosPress reproduces the high-quality patch features of the foundation model (Fig. 2), and can be used to produce specialised models, that have improved accuracy on particular tasks but retain these foundation model properties. In this work, we:
@@
  • show that CosPress can be used to train specialised models with improved performance on a particular vision task, but retain foundation model properties such as improved generalisability and out-of-distribution detection performance.
@@ 2 Related Work
Foundation models in computer vision follow the success of transformer-based foundation models in language, such as BERT [10], and encode images as vectors in latent space, where the distance between vectors describes the semantic similarity of the images. Two approaches have emerged for training these models: self-supervised learning [32] and contrastive language-image pretraining (CLIP) [34]. The CLIP models were among the first to show that by using a large Vision Transformer (ViT) architecture [11], and a large, diverse and high quality training dataset, a generalist vision model could be produced that achieves high performance across a range of applications [34]. The DINOv2 foundation models followed, and using a combined bootstrapping [14, 8] and masked patch prediction [57] approach to train foundation models with strong performance on image classification and segmentation tasks [32].
@@
Knowledge distillation is the process of transferring knowledge from a large model or model ensemble to a single smaller model. The earliest approaches aligned the output probability vectors of the student and teacher classifications using a Kullback–Leibler (KL) divergence loss [17]. There is a wide range of literature demonstrating how this approach can improve the performance of smaller Convolutional Neural Networks (CNNs) [45] and Vision Transformers (ViTs) [40, 50], by leveraging a strong teacher or one with different inductive biases to the student model. It has been shown that knowledge distillation is most effective when it is treated as a function matching problem [5], with the same inputs being provided to both the teacher and student model.
Feature distillation of Foundation models.
Feature distillation—where the output features of the teacher are used for training the student instead of the classification outputs—is less well studied for models without class outputs is less well studied. Generally speaking, feature distillation is used in combination with a knowledge distillation objective and often focuses on supervised models. The only method that we are aware of that considers distilling foundation models from scratch is Proteus [56], which finds that using only feature distillation objectives is important for preserving their properties. The components of Proteus—a student head to align output dimensions, MSE loss on class and patch tokens, and an iBOT [57] inspired masking objective—are a logical adaption of components in prior supervised feature distillation methods such as Masked Generative Distillation (MGD) [49], SRD [28] and $V_{k}D$ [29] to a ViT architecture with the aim of preserving both the local and global features of the teacher.
Frozen teacher $T$ Image $x$ Student $S_{\theta}$ Proteus loss $\mathcal{L}_{\text{proteus}}$ Student heads $g_{\theta},g_{\psi},g_{\nu}$ // Teacher head $h_{\phi}$ Student loss $\mathcal{L}_{\text{student}}$ Compression loss $\mathcal{L}_{\text{dim-red}}$ Cosine-similarity Preserving Compression Map //// CosPress (Ours) Proteus [56]
Figure 3: Foundation Model Feature distillation frameworks. In Proteus, student heads $g$ are used to map the outputs of the student network $S_{\theta}$ into the latent space of the teacher $T$, so that a MSE loss can be applied. In CosPress, a teacher head $h$ is trained to compress the teacher $T$ outputs into the student latent space, preserving the cosine similarity of image embeddings and allowing direct optimisation. The Proteus student head does not preserve cosine similarity, even when the projection matrices are forced to be right-orthogonal.
@@ 3 Methods
We are interested in the problem of training a student network to mimic the behaviour of a large, high quality teacher model using a ViT architecture, such as the DINOv2 foundation models. A key property of these models is that the cosine similarity between images and patch embeddings captures their semantic similarity, as shown by the use of this measure for zero-shot classification and identifying duplicate imagery [20, 32, 34]. However, larger ViT architectures have a larger output dimensionality, which prevents the embeddings produced by smaller student models being directly compared to a teacher model embedding.
Proteus [56] addresses this problem by introducing a student head $g:\mathbb{R}^{D_{S}}\to\mathbb{R}^{D_{T}}$ that maps the outputs of the student model into the latent space of the teacher model, allowing a MSE ($L_{2}$) loss to be applied. This student head contains a projection matrix, $\mathbf{W}\in\mathbb{R}^{D_{S}\times D_{T}}$, that maps the from the latent space of the student to the teacher, and is commonly discarded. Two problems with this approach are firstly, the projection can contain information for replicating the teacher network, and distort the outputs at the level of the indirectly optimised student model [29]. Secondly, it does not guarantee that the projection $\mathbf{W}$ will be faithful and preserve the cosine similarities between teacher model embeddings of different images, within the latent space of the student.
Prior work has found that requiring $\mathbf{W}$ to be right-orthogonal addresses the first problem [29]. This means that $\mathbf{W}\mathbf{W}^{\top}=\textbf{I}_{D_{S}}$ must be satisfied where $\textbf{I}_{D_{S}}\in\mathbb{R}^{D_{S}\times D_{S}}$ is the identity matrix of rank $D_{S}$. However, this implies that for any image $i$, we have
@@
$\frac{S_{\theta}^{c}(x_{i})\cdot{}S_{\theta}^{c}(x_{j})}{\|S_{\theta}^{c}(x_{i})\|\,\|S_{\theta}^{c}(x_{j})\|}$ $\approx\frac{T^{c}(x_{i})\mathbf{W}^{\top}\mathbf{W}T^{c}(x_{j})^{\top}}{\|T^{c}(x_{i})\mathbf{W}^{\top}\|\,\|\mathbf{W}T^{c}(x_{j})^{\top}\|}.$ (2)
This implies that $\mathbf{W}$ also needs to be left-orthogonal to address the second problem and ensure that cosine similarities are preserved. Asserting that with $\mathbf{W}^{\top}\mathbf{W}=\alpha\textbf{I}_{D_{T}}$ for some scalar $\alpha$ would yield the desired relationship
@@
Unfortunately, the projection $\mathbf{W}$ can only be approximately left-orthogonal, as $\mathbf{W}$ is only of rank $D_{S}$ and $\textbf{I}_{D_{T}}$ is of rank $D_{T}$. Nevertheless, it is found that approximate left-orthogonality implies right-orthogonality up to a scalar, as stated in the following Lemma 1.
@@
For any matrix $\mathbf{M}\in\mathbb{R}^{m\times d}$ with $m<d$ with rank $m$, we have the inequality $\|\mathbf{M}\mathbf{M}^{\top}-\frac{d}{m}\mathbf{I}_{m}\|_{F}\leq\|\mathbf{M}^{\top}\mathbf{M}-\mathbf{I}_{d}\|_{F}$. The converse inequality does not hold.
The proof of Lemma 1 stems from $\mathbf{M}$ having more columns than rows, which can be downsampled to form a square matrix, and is provided in Appendix A of the supporting information. While prior methods such as $V_{k}D$ introduce parametrisations that require $\mathbf{W}$ to be right-orthogonal [29], they do not also guarantee approximate left-orthogonality.
While Eq. 3 could be optimised directly using a Stochastic Neighbour Embedding (SNE) [18, 41] approach, our initial experiments found that this was less effective than Proteus due to this target having a complex loss surface with many local minima. Instead, we propose to use a teacher head, rather than a student head, to learn a function $h_{\phi}:\mathbb{R}^{D_{T}}\to\mathbb{R}^{D_{S}}$ that compresses the representation of the teacher into the latent space of the student while preserving cosine similarities
@@
which allows the student $S_{\theta}^{c}(x_{i})$ to be directly optimised against the compressed teacher representation $h_{\phi}(T^{c}(x_{i}))$. Learning the $h_{\phi}$ mapping is a tractable problem as described by the Johnson–Lindenstrauss (JL) Lemma, which states that such a mapping can be constructed with a margin of error that depends on the dimensionality of the target space and the size of the dataset of interest [13]. Further details are provided in Appendix A, and Fig. 3 highlights the differences between the Proteus and CosPress frameworks.
@@
In [56], the authors propose to minimize the MSE ($L_{2}$) loss between the outputs of the teacher and that of the student, once passed through a dimension-raising map called the student head $g:\mathbb{R}^{D_{S}}\to\mathbb{R}^{D_{T}}$. To achieve best performance, they use three student heads with different weights $\phi,\psi,\nu$ and minimise the $L_{2}$ loss separately on the class tokens, features (class and patch tokens), and on randomly masked tokens $\mathbf{X}^{M}$ similar to the MGD [49] approach. This leads to the following optimisation loss
@@
Secondly, the student network $S_{\theta}$ is trained to match the image of the teacher under the teacher head $h_{\phi}\circ T$ in the student loss term $\mathcal{L}_{\text{student}}$. The most effective way to fit this term is to freeze the teacher head $h_{\phi}$ gradients and train on both losses concurrently. Using a weighting scheme was observed to produce similar results (Appendix C).
@@
To build a loss function that will ensure the mapping $h_{\phi}$ satisfies Eq. 4 an SNE [18, 41] inspired approach is used. This involves defining a kernel to build distributions describing the similarity between vectors, allowing for embeddings in the high dimensional input space to be aligned with the low dimensional input space by minimising the KL divergence between these distributions.
@@
$k_{\tau}(y;z)$ $\propto\exp{}\left(\frac{y\cdot z}{\|y\|\,\|z\|}/\tau\right)$ (7)
with temperature hyperparameter $\tau$, so that Eq. 4 becomes $k_{\tau}(T^{c}(x_{i}),T^{c}(x_{j}))\approx k_{\tau}(h_{\phi}(T^{c}(x_{i})),h_{\phi}(T^{c}(x_{j})))$ for each $i,j$. Performing SNE, for a set of vectors in the input space $p_{i}\in\mathbf{p}$ and target space $q_{i}\in\mathbf{q}$ of size $N$, we construct the matrices $P^{\tau},Q^{\tau}$ that define the input and target distributions by
$P_{ij}^{\tau}=\frac{p_{j|i}+p_{i|j}}{2N},\quad p_{j|i}$ =$\frac{k_{\tau}(p_{i};p_{j})}{\sum_{i\not=k}k_{\tau}(p_{i};p_{k})}$ (8)
$Q_{ij}^{\tau}=\frac{q_{j|i}+q_{i|j}}{2N},\quad q_{j|i}$ =$\frac{k_{\tau}(q_{i};q_{j})}{\sum_{i\not=k}k_{\tau}(q_{i};q_{k})}$ (9)
where the first equation builds symmetric $P^{\tau},Q^{\tau}$ matrices, allowing for greater flexibility in the solution. If these $P^{\tau},Q^{\tau}$ matrices are equal, the cosine similarity between pairs of points in $\mathbf{p},\mathbf{q}$ will be equal and Eq. 4 will be satisfied. This can be achieved approximately by minimising the KL divergence $\left(D_{KL}\right)$ over $\bm{\tau}$, a vector of temperature values via
$L_{\text{KL}}(\mathbf{p},\mathbf{q})$ $=\frac{1}{|\bm{\tau}|}\sum_{\tau\in\bm{\tau}}D_{KL}(P^{\tau}\|Q^{\tau}).$ (10)
An ablation study on the best values of $\bm{\tau}$ is described in Tab. S20 in the supporting information.
@@
As done for the Proteus [56] student heads, the teacher head architecture in CosPress uses a LayerNorm [2] followed by a linear layer. This can be written as
@@
where $\bar{z}$ is the average of the $z$ vector elements, and the initialisation scheme sets the biases $\beta_{1},\beta_{2}$ to zero and the scaling $\gamma$ to one at the start of training. The linear map $\mathbf{W}^{\top}\in\mathbb{R}^{D_{T}\times D_{S}}$ is initialised using a random normal distribution as is standard, which is consistent with the mapping constructed in the JL Lemma (Appendix A). For the Proteus student heads $g$, $\mathbf{W}^{\top}$ is replaced by $\mathbf{W}$ and the dimension of the bias vectors are adjusted accordingly.
@@
$L_{\text{cosine}}(\mathbf{z},\mathbf{y})$ $=\frac{1}{n}\left(\sum_{i}1-\frac{z_{i}\cdot y_{i}}{\|z_{i}\|\,\|y_{i}\|}\right)$ (13)
where $\mathbf{z},\mathbf{y}$ are sets of input vectors of the same length $n$. Considering the teacher head $h_{\phi}$ learns to conserve cosine similarity, this is a natural choice for the student network and is found to result in better performance with CosPress than the $L_{2}$ loss (Appendix C).

### v3

@@ Abstract
Knowledge distillation methods compress models by training a student network using the classification outputs of a high quality teacher model, but can fail to effectively transfer the properties of computer vision foundation models from the teacher to the student. While it has been recently shown that feature distillation—where a teacher model’s output features are replicated instead—can reproduce performance for foundation models across numerous downstream tasks, they fall short in matching critical properties such as robustness and out-of-distribution (OOD) detection performance. This paper overcomes this shortcoming by introducing Cosine-similarity Preserving Compression (CosPress), a feature distillation technique that learns a mapping to compress the latent space of the teacher model into the smaller latent space of the student, by preserving the cosine similarities between image embeddings. This enables direct optimisation of the student network and produces a more faithful reproduction of the teacher’s properties. It is shown that distillation with CosPress on a variety of datasets, including ImageNet, produces more accurate models with greater performance on generalisability, robustness and OOD detection benchmarks, and that this technique provides a competitive pathway for training highly performant lightweight models on small datasets. Code is available at github.com/emannix/cospress.
1 Introduction
Deep learning computer vision approaches have become the standard for automating vision problems across a range of fields, from medical imaging (zhang2024challenges) to analysis of satellite imagery (bastani2023satlaspretrain) and detecting weapons in luggage (andriyanov2024intelligent). However, models trained with commonly used supervised learning approaches can have poor robustness (bai2021transformers; hendrycks2021natural) and struggle to detect out-of-distribution (OOD) data (yang2022openood; nguyen2015deep).
By leveraging large Vision Transformer (ViT) architectures and pretraining on large and diverse datasets, foundation models in computer vision comprise a significant step forward toward addressing these challenges, providing significantly improved generalisation ability and robustness in comparison to purely supervised approaches (oquab2023dinov2; radford2021learning). Large ViT models enjoy superior performance after pre-training (zhai2022scaling), and can be distilled to produce smaller models that are more practical for deployment. For example, smaller DINOv2 foundation models were distilled from their largest variant with 1.1 billion parameters by optimising the self-supervised training objective with a frozen teacher (oquab2023dinov2). This approach is not replicable, as it was conducted on the proprietary LVD-142M dataset and the teacher head weights were never publicly released.
@@ 1 Introduction
Nevertheless, knowledge distillation approaches (hinton2015distilling) that leverage the classification outputs of the DINOv2 models trained on a particular datasets can be used to train performant student models. However, it has been shown that fundamental properties of the foundation model need not transfer to the student models, impacting generalisation performance on downstream tasks (zhang2024accessingvisionfoundationmodels). Feature distillation approaches that use a Mean Squared Error (MSE) or $L_{2}$ loss and a student head to map the activations from the latent space of the student to the teacher are more effective in this respect. The recent Proteus (zhang2024accessingvisionfoundationmodels) approach has shown that it is possible to distill the DINOv2 models on ImageNet-1K and obtain comparable performance on downstream classification and segmentation tasks. However, sub-optimal results are still obtained for key robustness metrics, as well as generalisability and performance in dense tasks such as image segmentation. Most concerning, however, is that models distilled using Proteus do not faithfully reproduce the latent space of the teacher model, as shown by their severely reduced performance on out-of-distribution (OOD) detection tasks (Table˜1).
@@
This paper presents Cosine-similarity Preserving Compression (CosPress), a feature distillation approach that addresses such shortcomings by learning a teacher head mapping that compresses the latent space of the teacher model into the latent space of the student. CosPress achieves this by preserving cosine similarities between points in the teacher’s latent space and allows the student to be directly optimised. This significantly improves the faithfulness of the learned student on OOD detection (Table˜1) and robustness benchmarks in comparison to Proteus, while achieving competitive performance across all of the considered challenges—including classification accuracy, generalisation and semantic segmentation. CosPress reproduces the high-quality patch features of the foundation model (Fig.˜1), and can be used to produce specialised models, that have improved accuracy on particular tasks but retain these foundation model properties. In this work, we:
@@
  • show that CosPress can be used to train specialised models with improved performance on a particular vision task, while retaining foundation model properties such as improved generalisability and out-of-distribution detection performance.
@@ 2 Related Work
Foundation models in computer vision follow the success of transformer-based foundation models in language, such as BERT (devlin2018bert), and encode images as vectors in latent space, where the distance between vectors describes the semantic similarity of the images. Two approaches have emerged for training these models: self-supervised learning (oquab2023dinov2) and contrastive language-image pretraining (CLIP) (radford2021learning). The CLIP models were among the first to show that by using a large Vision Transformer (ViT) architecture (dosovitskiy2020image), and a large, diverse and high quality training dataset, a generalist vision model could be produced that achieves high performance across a range of applications (radford2021learning). The DINOv2 foundation models followed, and using a combined bootstrapping (grill2020bootstrap; caron2021emerging) and masked patch prediction (zhou2021ibot) approach to train foundation models with strong performance on image classification and segmentation tasks (oquab2023dinov2).
@@
Knowledge distillation is the process of transferring knowledge from a large model or model ensemble to a single smaller model. The earliest approaches aligned the output probability vectors of the student and teacher classifications using a Kullback–Leibler (KL) divergence loss (hinton2015distilling). There is a wide range of literature demonstrating how this approach can improve the performance of smaller Convolutional Neural Networks (CNNs) (wei2020circumventing) and Vision Transformers (ViTs) (touvron2021training; yang2024vitkd), by leveraging a strong teacher or one with different inductive biases to the student model. It has been shown that knowledge distillation is most effective when it is treated as a function matching problem (beyer2022knowledge), with the same inputs being provided to both the teacher and student model.
Feature distillation.
Feature distillation—where the output features of the teacher are used for training the student instead of the classification outputs—is less well studied for models without class outputs. Generally speaking, feature distillation is used in combination with a knowledge distillation objective and often focuses on supervised models. However, the Proteus approach (zhang2024accessingvisionfoundationmodels) demonstrated that using pure feature distillation objectives is important for preserving foundation model properties. The components of Proteus—a student head to align output dimensions, MSE loss on class and patch tokens, and an iBOT (zhou2021ibot) inspired masking objective—are a logical adaption of components in prior supervised feature distillation methods such as Masked Generative Distillation (MGD) (yang2022masked), SRD (miles2024understanding) and $V_{k}D$ (miles2024vkd) to a ViT architecture with the aim of preserving both the local and global features of the teacher.
Dimensionality reduction.
Feature distillation and the challenge of compressing latent spaces to train performant student models are closely related to the broader ideas of dimensionality reduction and minimum distortion embeddings (agrawal2021minimum). Stochastic Neighbor Embedding (SNE) (hinton2002stochastic; van2008visualizing) is a dimensionality reduction technique that projects high-dimensional embeddings into a low-dimensional space (typically two dimensions) while preserving local relationships. SNE acheives this by constructing a probability distribution over pairs of points in the original space and then optimising a corresponding set of points in the low-dimensional space to match this higher dimensional distribution as closely as possible. However, SNE does not learn an explicit mapping between the original and reduced spaces, only a lower dimensional representation. Other approaches, in contrast, explicitly learn projection functions, often with the goal of preserving local geometric structures such as distances or angles (saul2000introduction; he2003locality; gao2020robust; fischer2024sailing).
@@ 3 Methods
Frozen teacher $T$ Image $x$ Student $S_{\theta}$ Proteus loss $\mathcal{L}_{\text{proteus}}$ Student heads $g_{\theta},g_{\psi},g_{\nu}$ // Teacher head $h_{\phi}$ Student loss $\mathcal{L}_{\text{student}}$ Compression loss $\mathcal{L}_{\text{dim-red}}$ Cosine-similarity Preserving Compression Map //// CosPress Proteus
Figure 2: Feature distillation frameworks. In Proteus, student heads $g$ are used to map the outputs of the student network $S_{\theta}$ into the latent space of the teacher $T$, so that a MSE loss can be applied. In CosPress, a teacher head $h$ is trained to compress the teacher $T$ outputs into the student latent space, preserving the cosine similarity of image embeddings and allowing direct optimisation. The Proteus student head does not preserve cosine similarity, even when the projection matrices are forced to be right-orthogonal.
@@
We are interested in the problem of training a student network to mimic the behaviour of a large, high quality teacher model using a ViT architecture, such as the DINOv2 foundation models. A key property of these models is that the cosine similarity between images and patch embeddings captures their semantic similarity, as shown by the use of this measure for zero-shot classification and identifying duplicate imagery (jose2024dinov2; oquab2023dinov2; radford2021learning). However, larger ViT architectures have a larger output dimensionality, which prevents the embeddings produced by smaller student models being directly compared to a teacher model embedding.
Proteus (zhang2024accessingvisionfoundationmodels) addresses this problem by introducing a student head $g:\mathbb{R}^{D_{S}}\to\mathbb{R}^{D_{T}}$ that maps the outputs of the student model into the latent space of the teacher model, allowing a MSE ($L_{2}$) loss to be applied. This student head contains a projection matrix, $\mathbf{W}\in\mathbb{R}^{D_{S}\times D_{T}}$, that maps the from the latent space of the student to the teacher, and is commonly discarded. Two issues arise with this approach. First, the projection may encode information specific to replicating the teacher network, potentially distorting the outputs of the student model, which is only indirectly optimised (miles2024vkd). Second, there is no guarantee that the projection matrix $\mathbf{W}$ will be faithful in preserving cosine similarities between teacher embeddings of different images—similarities that reflect semantic relationships (jose2024dinov2)—within the student’s latent space.
Prior work has found that requiring $\mathbf{W}$ to be right-orthogonal addresses the first problem (miles2024vkd). Right-orthogonality means that $\mathbf{W}\mathbf{W}^{\top}=\textbf{I}_{D_{S}}$ must be satisfied where $\textbf{I}_{D_{S}}\in\mathbb{R}^{D_{S}\times D_{S}}$ is the identity matrix of rank $D_{S}$. However, this implies that for any image $i$, we have
@@
$\frac{S_{\theta}^{c}(x_{i})\cdot{}S_{\theta}^{c}(x_{j})}{\|S_{\theta}^{c}(x_{i})\|\,\|S_{\theta}^{c}(x_{j})\|}$ $\approx\frac{\left(T^{c}(x_{i})\mathbf{W}^{\top}\right)\cdot{}\left(\mathbf{W}\,T^{c}(x_{j})^{\top}\right)}{\|T^{c}(x_{i})\mathbf{W}^{\top}\|\,\|\mathbf{W}\,T^{c}(x_{j})^{\top}\|}.$ (2)
This implies that $\mathbf{W}$ also needs to be left-orthogonal to address the second problem and ensure that cosine similarities are preserved. Asserting left-orthogonality with $\mathbf{W}^{\top}\mathbf{W}=\alpha\textbf{I}_{D_{T}}$ for some scalar $\alpha$ would yield the desired relationship
@@
Unfortunately, the projection $\mathbf{W}$ can only be approximately left-orthogonal, as $\mathbf{W}$ is only of rank $D_{S}$, which means the product $\mathbf{W}^{\top}\mathbf{W}$ can only ever be of rank $D_{S}$. As $\textbf{I}_{D_{T}}$ is of rank $D_{T}>D_{S}$, then $\mathbf{W}^{\top}\mathbf{W}\not=\alpha\textbf{I}_{D_{T}}$. Consider the following definition.
Definition 1 (Approximately Orthogonal Matrix).
A matrix $\mathbf{M}\in\mathbb{R}^{m\times d}$ is said to be approximately orthogonal if
$\|\mathbf{M}\mathbf{M}^{\top}-\alpha\mathbf{I}_{m}\|_{F}<\varepsilon\quad\text{and}\quad\|\mathbf{M}^{\top}\mathbf{M}-\beta\mathbf{I}_{d}\|_{F}<\varepsilon,$
for sufficiently small $\varepsilon>0$ and real scalars $\alpha,\beta$. If only one of these conditions are satisfied the matrix is said to be approximately right or left orthogonal, as appropriate. This notion expands upon the idea of orthogonality, where a matrix $\mathbf{M}$ can only be orthogonal ($\varepsilon=0,\alpha=\beta=1$) if it is square ($m=d$).
@@
Let $\mathbf{M}\in\mathbb{R}^{m\times d}$ with $m<d$ and $\mathrm{rank}(\mathbf{M})=m$. Then
$\bigl\|\mathbf{M}\mathbf{M}^{\top}-\tfrac{d}{m}\mathbf{I}_{m}\bigr\|_{F}\;\leq\;\bigl\|\mathbf{M}^{\top}\mathbf{M}-\mathbf{I}_{d}\bigr\|_{F}.$
Moreover, the converse inequality does not generally hold.
This lemma shows that approximate right-orthogonality is sufficient for a matrix to also be approximately left-orthogonal, and therefore to be approximately orthogonal overall. Consequently, both conditions Eq.˜1 and Eq.˜3 are satisfied, ensuring that the mapping does not encode information while preserving the relationships between image embeddings. The proof of Lemma Lemma˜1 stems from $\mathbf{M}$ having more columns than rows, which can be downsampled to form a square matrix, and is provided in Appendix˜A of the supporting information. While prior methods such as $V_{k}D$ introduce parametrisations that require $\mathbf{W}$ to be right-orthogonal (miles2024vkd), they do not also guarantee approximate left-orthogonality.
While Eq.˜3 could be optimised directly using an SNE (hinton2002stochastic; van2008visualizing) inspired approach, our initial experiments found that this was less effective than Proteus due to this target having a complex loss surface with many local minima. Instead, it is proposed to use a teacher head, rather than a student head, to learn a function $h_{\phi}:\mathbb{R}^{D_{T}}\to\mathbb{R}^{D_{S}}$ that compresses the representation of the teacher into the latent space of the student while preserving cosine similarities
@@
which allows the student $S_{\theta}^{c}(x_{i})$ to be directly optimised against the compressed teacher representation $h_{\phi}(T^{c}(x_{i}))$. Learning the $h_{\phi}$ mapping is a tractable problem as described by the Johnson–Lindenstrauss (JL) Lemma, which states that such a mapping can be constructed with a margin of error that depends on the dimensionality of the target space and the size of the dataset of interest (freksen2021introduction). Further details are provided in Appendix˜A, and Fig.˜2 highlights the differences between the Proteus and CosPress frameworks.
@@
zhang2024accessingvisionfoundationmodels propose to minimize the MSE ($L_{2}$) loss between the outputs of the teacher and that of the student, when passed through a dimension-raising map called the student head $g:\mathbb{R}^{D_{S}}\to\mathbb{R}^{D_{T}}$. To achieve best performance, they use three student heads with different weights $\phi,\psi,\nu$ and minimise the $L_{2}$ loss separately on the class tokens, features (class and patch tokens), and on randomly masked tokens $\mathbf{X}^{M}$ similar to the MGD (yang2022masked) approach. This leads to the following optimisation loss
@@
Secondly, the student network $S_{\theta}$ is trained to match the image of the teacher under the teacher head $h_{\phi}\circ T$ in the student loss term $\mathcal{L}_{\text{student}}$. The most effective way to fit this term is to freeze the teacher head $h_{\phi}$ gradients and train on both losses concurrently as shown in Fig.˜2. Using a weighting scheme was observed to produce similar results (Appendix˜C).
@@
To build a loss function that will ensure the mapping $h_{\phi}$ satisfies Eq.˜4 an SNE (hinton2002stochastic; van2008visualizing) inspired approach is used. This involves defining a kernel to build distributions describing the similarity between vectors, allowing for embeddings in the high dimensional input space to be aligned with the low dimensional target space by minimising the KL divergence between these distributions.
@@
$k_{\tau}(y;z)$ $\propto\exp{}\left(\frac{y\cdot z}{\|y\|\,\|z\|}/\tau\right),$ (7)
with temperature hyperparameter $\tau$. As a result, Eq.˜4 becomes
$k_{\tau}\left(T^{c}(x_{i});T^{c}(x_{j})\right)\approx k_{\tau}\left(h_{\phi}(T^{c}(x_{i}));h_{\phi}(T^{c}(x_{j}))\right),$ (8)
for each $i,j$. Then, for a set of vectors in the input space $p_{i}\in\mathbf{p}$ and target space $q_{i}\in\mathbf{q}$ of size $N$, we construct the matrices $P^{\tau},Q^{\tau}$ that define the input and target distributions by
$P_{ij}^{\tau}=\frac{p_{j|i}+p_{i|j}}{2N},\quad p_{j|i}$ $=\frac{k_{\tau}(p_{i};p_{j})}{\sum_{i\not=k}k_{\tau}(p_{i};p_{k})},$ (9)
$Q_{ij}^{\tau}=\frac{q_{j|i}+q_{i|j}}{2N},\quad q_{j|i}$ $=\frac{k_{\tau}(q_{i};q_{j})}{\sum_{i\not=k}k_{\tau}(q_{i};q_{k})},$ (10)
where the first equation builds symmetric $P^{\tau},Q^{\tau}$ matrices, allowing for greater flexibility in the solution. If these $P^{\tau},Q^{\tau}$ matrices are equal, the cosine similarity between pairs of points in $\mathbf{p},\mathbf{q}$ will be equal and Eq.˜4 will be satisfied. This can be achieved approximately by minimising the KL divergence $\left(D_{\text{KL}}\right)$ over $\bm{\tau}$, a vector of temperature values via
$L_{\text{KL}}(\mathbf{p},\mathbf{q})$ $=\frac{1}{|\bm{\tau}|}\sum_{\tau\in\bm{\tau}}D_{\text{KL}}(P^{\tau}\|Q^{\tau}).$ (11)
An ablation study on the best values of $\bm{\tau}$ is described in Table˜S21 in the supporting information.
@@
The calculation of $L_{\text{KL}}(h_{\phi}(T^{c}(\mathbf{X})),T^{c}(\mathbf{X}))$ is the only term in the CosPress loss that is calculated between examples in a batch, and that will scale non-linearly with increasing batch size. All other terms in both CosPress and Proteus are computed within individual examples and scale linearly with batch size.
@@
As done for the Proteus (zhang2024accessingvisionfoundationmodels) student heads, the teacher head architecture in CosPress uses a LayerNorm (ba2016layer) followed by a linear layer. This can be written as
@@
where $\bar{z}$ is the average of the $z$ vector elements, and the initialisation scheme sets the biases $\beta_{1},\beta_{2}$ to zero and the scaling $\gamma$ to one at the start of training. The linear map $\mathbf{W}^{\top}\in\mathbb{R}^{D_{T}\times D_{S}}$ is initialised using a random normal distribution as is standard, which is consistent with the mapping constructed in the JL Lemma (Appendix˜A). For the Proteus student heads $g$, $\mathbf{W}^{\top}$ is replaced by $\mathbf{W}$ and the dimension of the bias vectors are adjusted accordingly.
@@
$L_{\text{cosine}}(\mathbf{z},\mathbf{y})$ $=\frac{1}{n}\left(\sum_{i}1-\frac{z_{i}\cdot y_{i}}{\|z_{i}\|\,\|y_{i}\|}\right),$ (14)
where $\mathbf{z},\mathbf{y}$ are sets of input vectors of the same length $n$. Considering the teacher head $h_{\phi}$ learns to conserve cosine similarity, this is a natural choice for the student network and is found to result in improved performance with CosPress than the $L_{2}$ loss (Appendix˜C).